### v1

# Taming Scalable Visual Tokenizer for Autoregressive Image Generation

Abstract
Figure 2: Effects of Distribution Gap on Codebook Usage. (a) T-SNE of the codebook (16,384 codebook size and 256 dimension) and sampled encoder features. (b) Codebook usage curve. The partial-update strategy adopted by VQGAN broadens the distribution gap between encoded features and non-activated codes, while those of IBQ based on all-codes updating are evenly mixed, maintaining a high codebook usage ($\sim$96%) throughout the training. Fully leveraging the codebook significantly improves the reconstruction quality.
Existing vector quantization (VQ) methods struggle with scalability, largely attributed to the instability of the codebook that undergoes partial updates during training. The codebook is prone to collapse as utilization decreases, due to the progressively widening distribution gap between non-activated codes and visual features. To solve the problem, we propose Index Backpropagation Quantization (IBQ), a new VQ method for the joint optimization of all codebook embeddings and the visual encoder. Applying a straight-through estimator on the one-hot categorical distribution between the encoded feature and codebook, all codes are differentiable and maintain a consistent latent space with the visual encoder. IBQ enables scalable training of visual tokenizers and, for the first time, achieves a large-scale codebook ($2^{18}$) with high dimension ($256$) and high utilization. Experiments on the standard ImageNet benchmark demonstrate the scalability and superiority of IBQ, achieving competitive results on both reconstruction ($1.00$ rFID) and autoregressive visual generation ($2.05$ gFID). The code and models are available at https://github.com/TencentARC/SEED-Voken.
Figure 1: Reconstruction and generation samples of IBQ. We show 1024 $\times$ 1024 reconstructed samples (top) and 256 $\times$ 256 generated samples (middle and bottom).
1 Introduction
Given the remarkable scalability and generalizability of autoregressive transformers [29, 2] in large language models [17, 26], recent efforts have been made to extend this success to visual generation [28, 6, 24, 25]. The common practice is to discretize images into tokens using vector quantization (VQ) before applying autoregressive transformers for discrete sequence modeling through next-token prediction.
Existing studies on autoregressive visual generation primarily aim to improve the reconstruction performance of visual tokenizers [6, 31, 32] or to refine the autoregressive modeling of transformers [25, 24, 16, 14]. The improvement of visual tokenizers is crucial, given that it can markedly amplify the potential performance of ensuing autoregressive models. Intuitively, scaling visual tokenizers by increasing the codebook size and embedding dimension could help bridge the gap between discrete and continuous representations, thereby mitigating the information loss associated with discrete tokens. However, it is noteworthy that current visual tokenizers [31, 24] have not demonstrated such scaling properties.
Empirical research [6, 24] has revealed that current VQ methods struggle with scalability due to the inherent tendency of the codebook to collapse. This arises because these methods only optimize limited selected codes during each backpropagation. Such a widely-adopted partial update strategy gradually broadens the distribution gap between non-activated codes and the visual encoder’s representation space, making the non-activated codes further less likely to be selected. As shown in Fig. 2, VQGAN [6] almost fails when scaling both codebook size (i.e., $16,384$) and embedding dimension (i.e., $256$) simultaneously. Only a small amount of codes share the same distribution with the visual encoder, and the codebook usage degrades from $68\%$ to $0.002\%$ after training one epoch.
To tackle the challenge, we introduce a new VQ method, namely, Index Backpropagation Quantization (IBQ). It globally updates the entire codebook in each backward process to ensure consistency with the distribution of the visual encoder. In such a way, all codes have the same probability of being selected, resulting in a high utilization of the codebook throughout the training process. Specifically, rather than directly applying the straight-through estimator [1] to the selected codes, we employ this reparameterization approach on the categorical distribution between visual features and all codebook embeddings, thereby rendering all codes differentiable. As shown in Fig. 2, the sampled visual features and the codebook embeddings from IBQ are evenly mixed. IBQ keeps a high codebook usage ($\sim 96\%$) throughout the training process. Fully leveraging the codebook effectively enhances the representation capacity, as demonstrated by the superior reconstruction of IBQ ($1.37$ rFID) compared to VQGAN ($3.98$ rFID).
Experiments on the standard ImageNet benchmark validate the scalability and superiority of our method in both reconstruction and generation. We conducted an in-depth study on the scaling law of the IBQ tokenizer from three aspects: codebook size, code dimension, and model size. We observed that the reconstruction performance or codebook usage significantly increases when scaling up tokenizer training. To our knowledge, IBQ is the pioneering work to train an extremely large codebook (i.e., $262,144$) with a relatively large code dimension (i.e., $256$). This achievement has resulted in state-of-the-art reconstruction performance, achieving a $1.00$ rFID. We employed the IBQ tokenizer in a vanilla autoregressive transformer for visual generation, training models of varying sizes from 300M to 2.1B, and achieved competitive results.
In summary, our contributions are threefold:
  • We propose a simple yet effective vector quantization method, dubbed Index Backpropagation Quantization (IBQ), for training scalable visual tokenizers.
  • We study the scaling properties of IBQ by increasing codebook size, code dimension, and model size. IBQ for the first time trains a super large codebook ($2^{18}$) with a large dimension ($256$) and high usage, achieving state-of-the-art reconstruction performance.
  • We present a family of vanilla autoregressive image generation models ranging from 300M to 2.1B, remarkably outperforming competing methods, e.g., LlamaGen [24], and Open-MAGVIT2 [16].
2 Related Work
2.1 Vector Quantization
At the core of visual tokenizers is vector quantization, which maps the visual signals into discrete tokens. VQ-VAE [28] proposes an encoder-quantizer-decoder structure and uses a learnable codebook as the discrete representation space. VQ-VAE2 [19] introduces multi-scale hierarchical VQ-VAE to enhance local features. VQGAN [6] further uses adversarial loss and perceptual loss for good perceptual quality. RQ-VAE [13] and DQ-VAE [9] improve VQGAN by residual quantization and dynamic quantization respectively. To improve codebook utilization for large-size codebooks, some works try to decrease code dimension [31, 24]. Following this observation, MAGVIT-v2 [32] reduces the code dimension to zero, and significantly enlarges the codebook size to $2^{18}$ with Lookup-Free Quantization. Instead of joint optimization of the model and codebook, VQGAN-LC [36] extends the codebook size to 100,000 by utilizing a frozen codebook with a trainable projector.
Existing VQ methods suffer from codebook collapse when scaling up tokenzier, and usually use a small codebook size or low code dimension, which limits the representational capacity of the codebook. In contrast, our proposed IBQ shows consistent improvements when scaling up code dimension, codebook size and model size.
2.2 Tokenized Visual Generation
Tokenizers map continuous visual signals into a discrete token sequence. For subsequent visual generation, there are two approaches, including non-autoregressive (NAR) and autoregressive (AR) generation. NAR [3, 32] usually adopts BERT-style transformers to predict masked tokens. For inference, these methods generate all tokens of an image simultaneously, and iteratively refine the generated images conditioned on the previous generation. In contrast, AR models perform next-token prediction in a raster-scan manner. VQGAN [6] adopts GPT2-medium architecture, while LlamaGen [24] employs Llama [26] for scalable image generation. VAR [25] extends “next-token prediction” to “next-scale prediction” and introduces adaptive normalization (AdaLN [18]) to improve visual generation quality. Open-MAGVIT2 [16] introduces asymmetric token factorization for super-large codebook learning. In this paper, we follow vanilla autoregressive visual generation, and adopt Llama [26] with AdaLn [18] for image generation.
Algorithm 1 Pseudocode of IBQ in a PyTorch-like style
def IBQ(z, codebook):
␣␣␣␣z:visualfeaturemap(B*h*w,D)
␣␣␣␣␣␣␣␣B:batchsize
␣␣␣␣␣␣␣␣h:heightoffeaturemap
␣␣␣␣␣␣␣␣w:widthoffeaturemap
␣␣␣␣␣␣␣␣D:featuredimension
␣␣␣␣codebook:(K,D)
␣␣␣␣␣␣␣␣K:codebooksize
␣␣␣␣␣␣␣␣D:codedimension
logits = mm(z, codebook.T) # (B * h * w, K)
Ind_soft = softmax(logits, dim=1) # (B * h * w, K)
_, indices = soft_one_hot.max(dim=1)
Ind_hard = onehot(indices) # (B * h * w, K)
Ind = Ind_hard - Ind_soft.detach() + Ind_soft
z_q = mm(Ind, codebook) # (B * h * w, D)
return z_q
 mm: matrix multiplication; onehot: transfer index into one-hot vector.
Figure 3: Gradient flow of different VQ methods. Naïve VQ and VQGAN/VQVAE only update the selected codes in each backward process. IBQ updates all codes simultaneously by transferring the gradients of soft one-hot categorical distribution to hard one-hot index.
3 Method
Our framework consists of two stages. The first stage is to learn a scalable visual tokenizer with high codebook utilization via Index Backpropagation Quantization. In the second stage, we use an autoregressive transformer for visual generation by next-token prediction.
3.1 Preliminary
Vector quantization (VQ) maps continuous visual signals into discrete tokens with a fixed-size codebook $\mathcal{C}\in\mathbb{R}^{K\times D}$, where $K$ is the codebook size and $D$ is the code dimension. Given an image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, VQ first utilizes an encoder to project the image into the feature map $\mathcal{Z}\in\mathbb{R}^{h\times w\times D}$, where $h=H/p$, $w=W/p$, and $p$ is the downsample ratio. Then the feature map is quantized into $\mathcal{Q}\in\mathbb{R}^{h\times w\times D}$ as discrete representations with the codebook. Finally, the decoder reconstructs the image given the quantized features.
For quantization, previous methods typically calculate the Euclidean distance between each visual feature $z\in\mathbb{R}^{D}$ and all codes of the codebook, and then select the closest code as its discrete representation [28, 6]. Since the $\operatorname*{arg\,min}$ operation in quantization is non-differentiable, they apply the straight-through estimator on selected codes to copy the gradients from the decoder, to optimize the encoder simultaneously. The quantization process is as follows:
$q$ $=\operatorname*{arg\,min}_{\mathcal{C}_{k}\in\mathcal{C}}||z-\mathcal{C}_{k}||\in\mathbb{R}^{D},$ (1)
$z_{q}$ $=z+\text{sg}[q-z],$ (2)
where sg[]delimited-[][$\cdot$][ ⋅ ] is stop-gradient operation.
The partial updating strategy (i.e., only selected codes are optimized) adopted by these methods progressively widens the distribution gap between visual features and non-activated codes. It incurs the instability during training due to the codebook collapse, which hampers the scalability of the visual tokenizer.
3.2 Index Backpropagation Quantization
Quantization. To ensure the consistent distribution between the codebook and encoded features through the training, we introduce an all-codes updating method, Index Backpropagation Quantization (IBQ). At the core of IBQ is passing gradients backward to all codes of the codebook, instead of the selected codes only. Algorithm 1 provides the pseudo-code of IBQ.
Specifically, we first perform dot product between the given visual feature $z$ and all code embeddings as logits and get probabilities (soft one-hot) by softmax function.
logits $=[z^{T}\mathcal{C}_{1},z^{T}\mathcal{C}_{2},\cdots,z^{T}\mathcal{C}_{K}]^{T}\in\mathbb{R}^{K},$ (3)
$\text{Ind}_{\text{soft}}$ $=\text{softmax}(\text{logits}),$ (4)
[…]
Then we copy the gradients of soft one-hot categorical distribution to hard one-hot index:
$\text{Ind}=\text{Ind}_{\text{hard}}-\text{sg}[\text{Ind}_{\text{soft}}]+\text{Ind}_{\text{soft}}.$ (6)
Given the index, the quantized feature is obtained by:
$z_{q}=\text{Ind}^{T}\mathcal{C}.$ (7)
In this way, we can pass the gradients to all codes of the codebook via index. By Index Backpropagation Quantization, the distribution of the whole codebook and encoded features remains consistent throughout completed training, thus gaining a high codebook utilization.
Training Losses. Similar to VQGAN [6], the tokenizer is optimized with a combination of losses:
$\mathcal{L}$ $=\mathcal{L}_{R}+\mathcal{L}_{Q}+\mathcal{L}_{P}+\mathcal{L}_{G}+\mathcal{L}_{E},$ (8)
where $\mathcal{L}_{R}$ is reconstruction loss of image pixels, $\mathcal{L}_{Q}$ is quantization loss between the selected code embeddings and encoded features, $\mathcal{L}_{P}$ is perceptual loss from LPIPS [34], $\mathcal{L}_{G}$ is adversarial loss with PatchGAN discriminator [10] to enhance the image quality, and $\mathcal{L}_{E}$ is entropy penalty to encourage codebook utilization.
We further introduce double quantization loss, to force the selected code embeddings and given encoded visual features towards each other.
${z_{q}}^{\prime}=$ $\text{Ind}_{\text{hard}}^{T}\cdot[\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{K}],$ (9)
$\begin{split}\mathcal{L}_{Q}=&||z_{q}-z||^{2}+\\ &||\text{sg}[z]-{z_{q}}^{\prime}||^{2}+\beta||z-\text{sg}[{z_{q}}^{\prime}]||^{2}.\end{split}$ (10)
Discussion with other VQ methods. Vector quantization (VQ) discretizes a continuous visual feature into a quantized code with a fix-size codebook. Naive VQ selects the code with the closest Euclidean distance to the given visual feature. Since the $\operatorname*{arg\,min}$ operation is non-differentiable, the gradient flow back to the encoder is truncated, and only the parameters of the decoder and the selected codes can be updated. Inspired by straight-through gradient estimator [1], VQ-VAE and VQGAN copy the gradients from the decoder to the encoder for gradient approximation. They also adopt quantization loss to optimize the selected codes towards the encoded features. Consequently, the VQ model can be trained in an end-to-end manner. Despite progress has been made, these VQ methods struggle to scale up tokenizers for better representation capacity.
As shown in Fig. 3, existing VQ methods simply optimize a limited number of codes towards encoded features within each backward process. It progressively broadens the distribution gap between the non-activated codes and the encoded features, eventually leading to codebook collapse. And this situation becomes more severe with the code dimension and codebook size increase. Rather than directly applying the straight-through estimator [1] on the selected codes, we employ this parameterization approach on the categorical distribution between visual features and all codebook embeddings to enable gradients backward to all codes. In such a way, the distribution between the whole codebook and the encoded features remains consistent throughout the training process. Consequently, IBQ achieves an extremely large codebook with high code dimension and utilization.
3.3 Autoregressive Transformer
After tokenization, the visual feature is quantized into discrete representations which are subsequently flattened in a raster-scan manner for visual generation. Given the discrete token index sequence $\mathcal{X}=\{x_{i}\}_{i=1}^{T}$, where $T=h^{\prime}\times w^{\prime}$, we leverage an autoregressive transformer to model the sequence dependency via next-token prediction. Specifically, the optimization process is to maximize the log-likelihood:
$p(x_{1},\cdots,x_{T}|c)=\prod_{t=1}^{T}p(x_{t}|x_{1},\cdots,x_{t-1},c),$ (11)
where $c$ is the condition such as class label.
Note that, since our focus is on the visual tokenizer, we adopt the vanilla architecture of autoregressive transformers akin to Llama [26] with AdaLN [18] for visual generation.
[…]
VQGAN [6] 2D 16 $\times$ 16 16 256 $\times$ 256 1,024 256 7.94 -- 44%
VQGAN [6] 2D 16 $\times$ 16 16 256 $\times$ 256 16,384 256 4.98 0.2843 5.9%
VQGAN [6] 2D 16 $\times$ 16 16 256 $\times$ 256 16,384 256 3.98 0.2873 5.3%
SD-VQGAN [20] 2D 16 $\times$ 16 16 256 $\times$ 256 16,384 8 5.15 -- --
MaskGIT [3] 2D 16 $\times$ 16 16 256 $\times$ 256 1,024 256 2.28 -- --
[…]
MaskBit [30] 2D 16 $\times$ 16 16 256 $\times$ 256 16,384 0 1.61 -- --
Open-MAGVIT2 [16] 2D 16 $\times$ 16 16 256 $\times$ 256 16,384 0 1.58 0.2261 100%
Open-MAGVIT2 [16] 2D 16 $\times$ 16 16 256 $\times$ 256 262,144 0 1.17 0.2038 100%
[…]
IBQ (Ours) 2D 16 $\times$ 16 16 256 $\times$ 256 262,144 256 1.00 0.2030 84%
Titok-L [33] 1D 32 -- 256 $\times$ 256 4,096 16 2.21 -- --
Titok-B [33] 1D 64 -- 256 $\times$ 256 4,096 16 1.70 -- --
Titok-S [33] 1D 128 -- 256 $\times$ 256 4,096 16 1.71 -- --
Table 1: Reconstruction performance of different tokenizers on $256\times 256$ ImageNet 50k validation set. reproduced VQGAN.
[…]
$3.98$ $0.2873$ $5.3\%$
$\checkmark$ 1.67 0.2340 98%
$\checkmark$ $\checkmark$ 1.55 0.2311 97%
$\checkmark$ $\checkmark$ $\checkmark$ 1.37 0.2235 96%
Table 2: Effectiveness of designed modules.
[…]
32 2.04 0.2408 92%
64 1.39 0.2281 69%
128 1.38 0.2255 77%
256 1.37 0.2235 96%
Table 3: Effect of codebook dim.
[…]
1,024 2.24 0.2580 99%
8,192 1.87 0.2437 98%
16,384 1.37 0.2235 96%
262,122 1.00 0.2030 84%
Table 4: Impact of codebook size.
[…]
size dim scale
LFQ 16,384 0 342M 1.58 0.2261 3.40 228.03
IBQ 16,384 256 342M 1.37 0.2235 2.88 254.73
Table 5: Performance comparison with LFQ.
[…]
1 1.80 0.2377 99%
2 1.55 0.2311 97%
4 1.37 0.2235 96%
Table 6: Benefit of larger model size.
Figure 4: Qualitative Reconstruction Comparison. We compare IBQ with LlamaGen and Open-MAGVIT2 tokenizer.
4 Experiment
Figure 5: Scaling up visual tokenizers (e.g., codebook size, code dimension and model size) improves visual soundness of reconstruction. Zoom in for a better view.
[…]
Diffusion ADM [5] 554M 10.94 101.0 0.69 0.63
CDM [8] -- 4.88 158.7 -- --
LDM-4 [20] 400M 3.60 247.7 -- --
DiT-XL/2 [18] 675M 2.27 278.2 0.83 0.57
VAR VAR-d16 [25] 310M 3.30 274.4 0.84 0.51
VAR-d20 [25] 600M 2.57 302.6 0.83 0.56
VAR-d24 [25] 1.0B 2.09 312.9 0.82 0.59
VAR-d30 [25] 2.0B 1.92 323.1 0.82 0.59
MAR MAR-B [14] 208M 2.31 281.7 0.82 0.57
MAR-L [14] 479M 1.78 296.0 0.81 0.60
MAR-H [14] 943M 1.55 303.7 0.81 0.62
AR VQGAN [6] 227M 18.65 80.4 0.78 0.26
VQGAN [6] 1.4B 15.78 74.3 -- --
VQGAN-re [6] 1.4B 5.20 280.3 -- --
ViT-VQGAN [31] 1.7B 4.17 175.1 -- --
ViT-VQGAN-re [31] 1.7B 3.04 227.4 -- --
RQTran. [13] 3.8B 7.55 134.0 -- --
RQTran.-re [13] 3.8B 3.80 323.7 -- --
LlamaGen-L [24] 343M 3.80 248.28 0.83 0.51
LlamaGen-XL [24] 775M 3.39 227.08 0.81 0.54
LlamaGen-XXL [24] 1.4B 3.09 253.61 0.83 0.53
LlamaGen-3B [24] 3.1B 3.06 279.72 0.84 0.53
LlamaGen-L [24] 343M 3.07 256.06 0.83 0.52
LlamaGen-XL [24] 775M 2.62 244.08 0.80 0.57
LlamaGen-XXL [24] 1.4B 2.34 253.90 0.80 0.59
LlamaGen-3B [24] 3.1B 2.18 263.33 0.81 0.58
Open-MAGVIT2-B [16] 343M 3.08 258.26 0.85 0.51
Open-MAGVIT2-L [16] 804M 2.51 271.70 0.84 0.54
Open-MAGVIT2-XL [16] 1.5B 2.33 271.77 0.84 0.54
AR IBQ-B (Ours) 342M 2.88 254.73 0.84 0.51
IBQ-L (Ours) 649M 2.45 267.48 0.83 0.52
IBQ-XL (Ours) 1.1B 2.14 278.99 0.83 0.56
IBQ-XXL (Ours) 2.1B 2.05 286.73 0.83 0.57
Table 7: Class-conditional generation on $256\times 256$ ImageNet. * specifies the generated images are $384\times 384$ and are resized to 256×256 for evaluation. The evaluation protocol and implementation are the same with ADM.
4.1 Datasets and Metrics
The training of the visual tokenizer and autoregressive transformer are both on $256\times 256$ ImageNet [4]. For visual reconstruction, the reconstruction-FID, denoted as rFID [7], codebook utilization, and LPIPS [34] on ImageNet 50k validation set are adopted to measure the quality of reconstructed images. For visual generation, we measure the quality of image generation by the prevalent metrics FID, IS [21] and Precision/Recall [12].
4.2 Implementations Details
Visual Reconstruction Setup.
We adopt the same model architecture proposed in VQGAN [6]. The visual tokenizer is trained with the following settings: an initial $41e-$ 4 learning rate with $0.01$ multi-step decay mechanism, an Adam Optimizer [11] with $\beta_{1}=0.5$, $\beta_{2}=0.9$, a total $256$ batch size with $330$ epochs, a combination of reconstruction, GAN [10], perceptual [34], commitment [6], entropy [32], double quantization losses, and LeCAM regularization [27] for training stability. When not specified, we use 16,384 codebook size, 256 code dimension, and 4 ResBlocks as our default tokenizer setting.
Visual Generation Setup.
It is anticipated that the plain autoregressive visual generation method can well exhibit the effectiveness of our visual tokenizer design. Specifically, we adopt Llama-based architecture where the RoPE [23], SwiGLU [22], and RMSNorm [35] techniques are incorporated. We further introduce AdaLN [18] for better visual synthesis quality. The class embedding serves as the start token and the conditional of AdaLN simultaneously. IBQ with width $w$, depth $d$ and head $h$ follows the scaling rules proposed in [24, 25], which is formulated as:
$w=64d,\quad h=d.$ (12)
All models are trained with similar settings: a base learning rate of $41e-$ 4 per $256$ batch size, an AdamW optimizer [15] with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay = $25e-$ 2, a total $768$ batch size and 300 $\sim$ 450 training epochs corresponding the model size, gradient clipping of $1.0$, and $0.1$ dropout rate for input embedding, FFN module and conditional embedding.
4.3 Main Results
Visual Reconstruction.
Tab. 1 shows the quantitative reconstruction comparison between IBQ and prevalent visual tokenizers. It can be seen that a significant drop in codebook usage on existing VQ methods when scaling codebook size (e.g., VQGAN [6] has a 44% usage for 1024 codebook size, while a 5.9% usage for 16,384 codebook size.), and code dimension (e.g., LlamaGen [24] has a 97% usage for 8-dimension codes, while a 0.29% usage for 256-dimension codes.) Therefore, the actual representational capacity is limited by the codebook collapse.
In contrast, our joint optimization of all codebook embedding and the visual encoder ensures the consistent distribution between them, facilitating the stable training of an large codebook size and embedding visual tokenizer with high utilization. Specifically, IBQ with 16,384 codebook size and 256 code dimension achieves 1.37 rFID, exceeding other VQ methods under the same downsampling rate and codebook size. By increasing the codebook size to 262,144, IBQ outperforms Open-MAGVIT2 [16] and achieves state-of-the-art reconstruction performance (1.00 rFID). We also perform a qualitative comparison with several representative VQ methods in Fig. 4. IBQ exhibits better visual soundness on complex scenarios such as faces and characters.
Visual Generation.
In Tab. 7, we compare IBQ with other generative models, including Diffusion models, AR models, and variants of AR models (VAR [25] and MAR [14]) on class-conditional image generation task. With the powerful visual tokenizers of IBQ, our models show consistent improvements when scaling up the model size (from 300M to 2.1B), and outperform all previous vanilla autoregressive models at different scales of model size. Moreover, IBQ outperforms the diffusion-based model DiT [18], and achieves comparable results with the variants of AR models. These AR model variants focus on the architecture designs of transformers in the second stage, while our work is devoted to better visual tokenizers in the first stage. Therefore, we believe that with our stronger tokenizers, the AR models and their variants can be boosted further.
4.4 Scaling Up IBQ
[…]
IBQ shows promising scaling capacity in three aspects: 1) Codebook Size: As shown in Tab. 6, there is a significant improvement in reconstruction quality as the codebook size enlarges from 1024 to 16,384. Furthermore, IBQ enables high codebook utilization and consistent gain in visual soundness even training with 262,144 codes. 2) Model Size: Tab. 6 reveals that by extending the number of ResBlock both in the encoder and the decoder, the boost in reconstruction performance can be guaranteed. 3) Code Dimension: interestingly, we observe significant codebook usage increasements when scaling up code dimension. We assume that low-dimensional codes are less discriminative where similar ones tend to be clustered. It indicates that the representative codes are more likely to be selected under our global updating strategy. In contrast, codes with high-dimensional embedding are highly informative since they are mutually sparse in the representation space. Therefore, these codes can be evenly selected during training, which ensures high utilization with better performance. With the above factors, we realize a super large codebook of 262,144 codebook size and 256 dimensions with high codebook usage (84%), achieving the state-of-the-art reconstruction performance (1.00 rFID). To better illustrate the scaling properties, we also provide visualizations in Fig. 5.
4.5 Ablation Studies
Key Designs.
[…]
Comparison with LFQ.
For fair comparisons, we adopt LFQ [16] with 16,384 codes and replace its asymmetric token factorization with our vanilla transformer architecture. We compare with LFQ on both reconstruction and generation in Tab. 6, and our proposed IBQ performs better, which demonstrates increasing code dimension can improve the reconstruction ability of the visual tokenizer and further boost the visual generation.
5 Conclusion
In this paper, we identify the bottleneck in scaling tokenizers arising from the partial-update strategy in current VQ methods, which progressively broadens the distribution gap between encoded features and non-activated codes, ultimately leading to codebook collapse. To tackle this challenge, we propose a simple yet effective vector quantization method, coined as Index Backpropagation Quantization (IBQ), for scalable tokenizer training, which updates all codes by applying the straight-through estimator on the categorical distribution between visual features and all codebook embeddings, thereby maintaining consistent distribution between the entire codebook and encoded features. Experiments on ImageNet demonstrate that IBQ enables a high-utilization, large-scale visual tokenizer with improved performance in both reconstruction (1.00 rFID) and generation (2.05 gFID), confirming the scalability and effectiveness of our method.
Appendix
Appendix A Autoregressive Model Configurations
We show the detailed autoregressive model configurations in Tab. 8. We scale up the AR models from 342M to 2.1B following the scaling rules proposed in VAR [25].
Appendix B Comparison with Soft Vector Quantization
To comprehensively illustrate the rationality of our IBQ, we compare it with another global update method, Soft Vector Quantization dubbed as Soft VQ. During training, it adopts the weighted average of all code embeddings as the quantized feature $v_{q}$ and incorporates a cosine decay schedule of the temperature ranging from $0.9$ to $1e$-$6$ for one-hot vector approximation. As for inference, it switches back to the original VQGAN way, which selects the code with the highest probability for hard quantization.
As shown in Tab. 9, Soft VQ is far behind IBQ in both reconstruction quality and codebook usage. In the experiments, we observe that the training process of Soft VQ corrupts within a few epochs ($<10$). This may stem from the unstable adversarial training where the adaptive weight of the GAN loss appears enormous and ends up with NAN. In addition, the soft-to-hard manner for one-hot vector approximation brings more difficulty in optimization and incurs inconsistency of quantization between training and inference, as demonstrated by a significant reconstruction quality drop ($16.17\text{rFID}\rightarrow 233.17\text{rFID}$).
Moreover, we provide an in-depth investigation by visualizing the distribution between the codebook and encoded features of Soft VQ. As shown in Fig. 6, although all-code updating strategy is enabled, the inappropriate quantization process tends to cluster codes mistakenly which subsequently suffers from low codebook usage (2.5%). We speculate that the force of the weighted average of code embeddings toward the encoded feature will smooth the codebook representation and result in similar and less informative code embeddings. In contrast, IBQ adopts hard quantization with index backpropagation. The hard quantization only involves the selected codes toward the encoded features for discriminative representation, thus ensuring precise quantization, while index backpropagation performs joint optimization of all codebook and the visual encoder to achieve consistent distribution. Considering the factors above, our proposed IBQ shows dominance in both reconstruction quality and codebook utilization.
Appendix C Additional Visualizations
We provide more qualitative reconstruction and generation samples in Fig. 7 and Fig. 8, respectively.
Appendix D Limitation and Future Work
In this work, we mainly focus on visual tokenizers through scalable training, and adopt vanilla autoregressive transformers for visual generation. We believe that combining our powerful visual tokenizers with advanced AR models is a promising way to improve autoregressive visual generation. In addition, our models are trained on ImageNet [4], limiting the generalization ability. Pretraining tokenizers on more data and training AR models on text-image pairs would be helpful. We leave these for future work.
Figure 6: Distribution Gap. The T-SNE results of the codebook (16,384 codebook size and 256 dimension) and sampled encoded features.
Model Parameters Width $w$ Head $h$ Depth $d$
IBQ-B 342M 16 16 1024
IBQ-L 649M 20 20 1280
IBQ-XL 1.1B 24 24 1536
IBQ-XXL 2.1B 30 30 1920
Table 8: Model sizes and architecture configurations of IBQ.
[…]
Soft VQ Corrupted Soft Soft 16.17 2.5%
Soft VQ Corrupted Soft Hard 233.17 2.5%
IBQ (Ours) Stable Hard Hard 4.03 99%
IBQ (Ours) Stable Hard Hard 1.37 96%
Table 9: Comparison with Soft Vector Quantization. Soft VQ training corrupts after a few epochs. When adopting hard quantization for inference, there is a significant drop in rFID. denotes IBQ with the same training epochs as Soft VQ.
Figure 7: The upper part illustrates the IBQ tokenizer tested at 1024 $\times$ 1024 Unsplash. While the second part showcases the IBQ tokenizer tested at 256 $\times$ 256 Imagenet. (a) indicates the original images and (b) signifies the reconstructions.
Figure 8: We showcases the 256 $\times$ 256 class conditional generation samples on Imagenet.

### v2

@@
-Taming Scalable Visual Tokenizer for Autoregressive Image Generation
+Scalable Image Tokenization with Index Backpropagation Quantization
 Abstract
@@
 (b) Codebook usage curve.
-The partial-update strategy adopted by VQGAN broadens the distribution gap between encoded features and non-activated codes, while those of IBQ based on all-codes updating are evenly mixed, maintaining a high codebook usage ($\sim$96%) throughout the training.
+The partial-update strategy adopted by VQGAN broadens the distribution gap between encoder features and non-activated codes, while those of IBQ based on all-codes updating are evenly mixed, maintaining a high codebook usage ($\sim$96%) throughout the training.
 Fully leveraging the codebook significantly improves the reconstruction quality.
@@
 IBQ enables scalable training of visual tokenizers and, for the first time, achieves a large-scale codebook ($2^{18}$) with high dimension ($256$) and high utilization.
-Experiments on the standard ImageNet benchmark demonstrate the scalability and superiority of IBQ, achieving competitive results on both reconstruction ($1.00$ rFID) and autoregressive visual generation ($2.05$ gFID).
+Experiments on the standard ImageNet benchmark demonstrate the scalability and superiority of IBQ, achieving competitive results on reconstruction and the application of autoregressive visual generation.
 The code and models are available at https://github.com/TencentARC/SEED-Voken.
@@
 1 Introduction
-Given the remarkable scalability and generalizability of autoregressive transformers [29, 2] in large language models [17, 26], recent efforts have been made to extend this success to visual generation [28, 6, 24, 25].
-The common practice is to discretize images into tokens using vector quantization (VQ) before applying autoregressive transformers for discrete sequence modeling through next-token prediction.
-Existing studies on autoregressive visual generation primarily aim to improve the reconstruction performance of visual tokenizers [6, 31, 32] or to refine the autoregressive modeling of transformers [25, 24, 16, 14].
-The improvement of visual tokenizers is crucial, given that it can markedly amplify the potential performance of ensuing autoregressive models.
-Intuitively, scaling visual tokenizers by increasing the codebook size and embedding dimension could help bridge the gap between discrete and continuous representations, thereby mitigating the information loss associated with discrete tokens.
+Discrete tokenizer plays a pivotal role in processing complex data across various modalities, such as text [6, 26, 40], images [42, 11, 3], and audio [2, 5].
+By transforming raw data into discrete tokens, models can effectively handle diverse data types within a unified framework [43, 6], simplifying the integration of multimodal information and facilitating native large multimodal models [38, 44].
+In the image domain, pioneering works like VQGAN [11] employ vector quantization (VQ) to learn visual tokenizers, enabling effective data compression and reconstruction.
+VQ tokenizers are deemed as the key component in applications such as autoregressive image generation [42, 11, 37, 39] and representation learning [3, 22].
+However, a notable challenge with VQ-based methods is the information loss during quantization, leading to inferior reconstruction performance compared to continuous representation models like Variational Autoencoders (VAEs) [17, 31].
+Intuitively, scaling visual tokenizers by increasing the codebook size and embedding dimension could help mitigate the information loss associated with discrete tokens, thereby bridging the gap between discrete and continuous representations.
 However, it is noteworthy that current visual tokenizers [46, 37] have not demonstrated such scaling properties.
 Empirical research [11, 37] has revealed that current VQ methods struggle with scalability due to the inherent tendency of the codebook to collapse.
-This arises because these methods only optimize limited selected codes during each backpropagation.
+This arises because these methods only optimize a limited number of the selected codes during each backpropagation.
 Such a widely-adopted partial update strategy gradually broadens the distribution gap between non-activated codes and the visual encoder’s representation space, making the non-activated codes further less likely to be selected.
@@
 IBQ keeps a high codebook usage ($\sim 96\%$) throughout the training process.
-Fully leveraging the codebook effectively enhances the representation capacity, as demonstrated by the superior reconstruction of IBQ ($1.37$ rFID) compared to VQGAN ($3.98$ rFID).
-Experiments on the standard ImageNet benchmark validate the scalability and superiority of our method in both reconstruction and generation.
-We conducted an in-depth study on the scaling law of the IBQ tokenizer from three aspects: codebook size, code dimension, and model size.
-We observed that the reconstruction performance or codebook usage significantly increases when scaling up tokenizer training.
+Fully utilizing the codebook effectively enhances the representation capacity, as demonstrated by the superior reconstruction of IBQ ($1.37$ rFID) compared to VQGAN ($3.98$ rFID).
+We conduct a comprehensive study on the scaling behavior of IBQ tokenizers along three axes: codebook size, code dimension, and model size.
+We observe significant gains in reconstruction quality or codebook usage when scaling up tokenizers.
 To our knowledge, IBQ is the pioneering work to train an extremely large codebook (i.e., $262,144$) with a relatively large code dimension (i.e., $256$).
-This achievement has resulted in state-of-the-art reconstruction performance, achieving a $1.00$ rFID.
-We employed the IBQ tokenizer in a vanilla autoregressive transformer for visual generation, training models of varying sizes from 300M to 2.1B, and achieved competitive results.
+This achievement leads to state-of-the-art reconstruction quality, reaching an rFID of $1.00$.
+We further demonstrate the effectiveness of IBQ tokenizers in autoregressive image generation by integrating them with vanilla transformers of varying scales, ranging from 300M to 2.1B parameters, achieving competitive performance.
+Although IBQ is also compatible with other advanced autoregressive models, the paper does not focus on them, leaving them for future research.
 In summary, our contributions are threefold:
@@
 IBQ for the first time trains a super large codebook ($2^{18}$) with a large dimension ($256$) and high usage, achieving state-of-the-art reconstruction performance.
-  • We present a family of vanilla autoregressive image generation models ranging from 300M to 2.1B, remarkably outperforming competing methods, e.g., LlamaGen [24], and Open-MAGVIT2 [16].
+  • We validate the effectiveness of IBQ tokenizers in visual generation by equipping them with vanilla autoregressive models ranging from 300M to 2.1B, remarkably outperforming competing methods, e.g., LlamaGen [37], and Open-MAGVIT2 [25].
 2 Related Work
@@
 At the core of visual tokenizers is vector quantization, which maps the visual signals into discrete tokens.
-VQ-VAE [28] proposes an encoder-quantizer-decoder structure and uses a learnable codebook as the discrete representation space.
+VQ-VAE [42] proposes an encoder-quantizer-decoder structure with a learnable codebook as the discrete representation space.
 VQ-VAE2 [30] introduces multi-scale hierarchical VQ-VAE to enhance local features.
 VQGAN [11] further uses adversarial loss and perceptual loss for good perceptual quality.
-RQ-VAE [13] and DQ-VAE [9] improve VQGAN by residual quantization and dynamic quantization respectively.
+RQ-VAE [21] and DQ-VAE [15] improve VQGAN by residual quantization and dynamic quantization, respectively.
 To improve codebook utilization for large-size codebooks, some works try to decrease code dimension [46, 37].
-Following this observation, MAGVIT-v2 [32] reduces the code dimension to zero, and significantly enlarges the codebook size to $2^{18}$ with Lookup-Free Quantization.
-Instead of joint optimization of the model and codebook, VQGAN-LC [36] extends the codebook size to 100,000 by utilizing a frozen codebook with a trainable projector.
-Existing VQ methods suffer from codebook collapse when scaling up tokenzier, and usually use a small codebook size or low code dimension, which limits the representational capacity of the codebook.
-In contrast, our proposed IBQ shows consistent improvements when scaling up code dimension, codebook size and model size.
+Following this observation, MAGVIT-v2 [47] reduces the code dimension to zero, and expands the codebook size to $2^{18}$ with Lookup-Free Quantization.
+Instead of joint optimization of the model and codebook, VQGAN-LC [52] extends the codebook size to 100,000 using a frozen codebook with a trainable projector.
+However, it introduces a bottleneck that constrains the tokenizer capacity.
+Existing VQ methods suffer from codebook collapse when scaling up tokenizers and typically use small-size codebooks or low-dimensional code embeddings, limiting the representational capacity.
+In contrast, our proposed IBQ shows consistent improvements when scaling up codebook size, code dimension and model size.
 2.2 Tokenized Visual Generation
@@
 VQGAN [11] adopts GPT2-medium architecture, while LlamaGen [37] employs Llama [40] for scalable image generation.
-VAR [25] extends “next-token prediction” to “next-scale prediction” and introduces adaptive normalization (AdaLN [18]) to improve visual generation quality.
-Open-MAGVIT2 [16] introduces asymmetric token factorization for super-large codebook learning.
-In this paper, we follow vanilla autoregressive visual generation, and adopt Llama [26] with AdaLn [18] for image generation.
+VAR [39] extends “next-token prediction” to “next-scale prediction” and introduces adaptive normalization (AdaLN [28]) to improve generation quality.
+Open-MAGVIT2 [25] proposes asymmetric token factorization for super-large codebook learning.
+In this paper, we adopt vanilla autoregressive models to validate the effectiveness of IBQ tokenizers in visual generation, excluding masked modeling or multi-scale structures for simplicity.
+Notably, IBQ is compatible with advanced generative models, which can further unlock the tokenizer potential.
+We leave this exploration for future work.
 Algorithm 1 Pseudocode of IBQ in a PyTorch-like style
@@
 Figure 3: Gradient flow of different VQ methods.
-Naïve VQ and VQGAN/VQVAE only update the selected codes in each backward process.
+VQGAN/VQVAE only update the selected codes in each backward process.
 IBQ updates all codes simultaneously by transferring the gradients of soft one-hot categorical distribution to hard one-hot index.
 3 Method
-Our framework consists of two stages.
-The first stage is to learn a scalable visual tokenizer with high codebook utilization via Index Backpropagation Quantization.
-In the second stage, we use an autoregressive transformer for visual generation by next-token prediction.
-3.1 Preliminary
+3.1 Preliminary: Vector Quantization
 Vector quantization (VQ) maps continuous visual signals into discrete tokens with a fixed-size codebook $\mathcal{C}\in\mathbb{R}^{K\times D}$, where $K$ is the codebook size and $D$ is the code dimension.
 Given an image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, VQ first utilizes an encoder to project the image into the feature map $\mathcal{Z}\in\mathbb{R}^{h\times w\times D}$, where $h=H/p$, $w=W/p$, and $p$ is the downsample ratio.
-Then the feature map is quantized into $\mathcal{Q}\in\mathbb{R}^{h\times w\times D}$ as discrete representations with the codebook.
+The feature map is then quantized into $\mathcal{Q}\in\mathbb{R}^{h\times w\times D}$ discrete representations using the codebook.
 Finally, the decoder reconstructs the image given the quantized features.
-For quantization, previous methods typically calculate the Euclidean distance between each visual feature $z\in\mathbb{R}^{D}$ and all codes of the codebook, and then select the closest code as its discrete representation [28, 6].
-Since the $\operatorname*{arg\,min}$ operation in quantization is non-differentiable, they apply the straight-through estimator on selected codes to copy the gradients from the decoder, to optimize the encoder simultaneously.
-The quantization process is as follows:
+Previous methods quantize each visual feature $z\in\mathbb{R}^{D}$ by selecting the nearest code from the codebook based on Euclidean distance [42, 11].
+Since the $\operatorname*{arg\,min}$ operation in quantization is non-differentiable, they apply the straight-through estimator on the selected codes to copy the gradients from the decoder to the encoder, to optimize the encoder and decoder simultaneously.
+The quantization process can be formulated as:
 $q$ $=\operatorname*{arg\,min}_{\mathcal{C}_{k}\in\mathcal{C}}||z-\mathcal{C}_{k}||\in\mathbb{R}^{D},$ (1)
@@
 where sg[]delimited-[][$\cdot$][ ⋅ ] is stop-gradient operation.
-The partial updating strategy (i.e., only selected codes are optimized) adopted by these methods progressively widens the distribution gap between visual features and non-activated codes.
+This partial updating strategy (i.e., only selected codes are optimized) adopted by these methods progressively widens the distribution gap between visual features and non-activated codes.
 It incurs the instability during training due to the codebook collapse, which hampers the scalability of the visual tokenizer.
@@
 To ensure the consistent distribution between the codebook and encoded features through the training, we introduce an all-codes updating method, Index Backpropagation Quantization (IBQ).
-At the core of IBQ is passing gradients backward to all codes of the codebook, instead of the selected codes only.
+The core idea of IBQ is to pass gradients to all codebook embeddings, rather than the selected ones only.
 Algorithm 1 provides the pseudo-code of IBQ.
@@
 […]
-Then we copy the gradients of soft one-hot categorical distribution to hard one-hot index:
+We then copy the gradients from soft one-hot categorical distribution to hard one-hot index:
 $\text{Ind}=\text{Ind}_{\text{hard}}-\text{sg}[\text{Ind}_{\text{soft}}]+\text{Ind}_{\text{soft}}.$ (6)
-Given the index, the quantized feature is obtained by:
+Given the index, the quantized feature can be computed as:
 $z_{q}=\text{Ind}^{T}\mathcal{C}.$ (7)
@@
 $\mathcal{L}$ $=\mathcal{L}_{R}+\mathcal{L}_{Q}+\mathcal{L}_{P}+\mathcal{L}_{G}+\mathcal{L}_{E},$ (8)
-where $\mathcal{L}_{R}$ is reconstruction loss of image pixels, $\mathcal{L}_{Q}$ is quantization loss between the selected code embeddings and encoded features, $\mathcal{L}_{P}$ is perceptual loss from LPIPS [34], $\mathcal{L}_{G}$ is adversarial loss with PatchGAN discriminator [10] to enhance the image quality, and $\mathcal{L}_{E}$ is entropy penalty to encourage codebook utilization.
+where $\mathcal{L}_{R}$ is reconstruction loss of image pixels, $\mathcal{L}_{Q}$ is quantization loss between the selected code embeddings and encoded features, $\mathcal{L}_{P}$ is perceptual loss from LPIPS [50], $\mathcal{L}_{G}$ is adversarial loss with PatchGAN discriminator [16] to enhance the image quality, and $\mathcal{L}_{E}$ is entropy penalty to encourage codebook utilization [47].
+To better explain how IBQ keeps the consistent distribution between the encoder features and the whole codebook, we provide a gradient analysis.
+Considering the quantization loss $\mathcal{L}_{Q}=||z-z_{q}||^{2}$,
+$\frac{\partial\mathcal{L}_{Q}}{\partial\mathcal{C}_{k}}$ $=-2\text{Ind}_{k}(z-z_{q})=-2p_{k}(z-z_{q}),$ (9)
+$\ p_{k}$ $=\frac{\text{exp}(z^{T}\mathcal{C}_{k})}{\sum_{j=1}^{K}\text{exp}(z^{T}\mathcal{C}_{j})}.$ (10)
+The softmax probabilities $p_{k}$ ensure that each $\mathcal{C}_{k}$ is updated based on its similarity to the encoder feature $z$, and $z-z_{q}$ shifts $\mathcal{C}_{k}$ toward dominant regions of the feature distribution $P_{Z}(z)$.
+Random batch sampling covers the whole encoder latent space, gradually aligning the entire codebook $\mathcal{C}$ with the distribution $P_{Z}(z)$ of encoder features over time.
 We further introduce double quantization loss, to force the selected code embeddings and given encoded visual features towards each other.
-${z_{q}}^{\prime}=$ $\text{Ind}_{\text{hard}}^{T}\cdot[\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{K}],$ (9)
+${z_{q}}^{\prime}=$ $\text{Ind}_{\text{hard}}^{T}\mathcal{C},$ (11)
 $\begin{split}\mathcal{L}_{Q}=&||z_{q}-z||^{2}+\\ &||\text{sg}[z]-{z_{q}}^{\prime}||^{2}+\beta||z-\text{sg}[{z_{q}}^{\prime}]||^{2}.\end{split}$ (12)
 Discussion with other VQ methods.
-Vector quantization (VQ) discretizes a continuous visual feature into a quantized code with a fix-size codebook.
-Naive VQ selects the code with the closest Euclidean distance to the given visual feature.
-Since the $\operatorname*{arg\,min}$ operation is non-differentiable, the gradient flow back to the encoder is truncated, and only the parameters of the decoder and the selected codes can be updated.
-Inspired by straight-through gradient estimator [1], VQ-VAE and VQGAN copy the gradients from the decoder to the encoder for gradient approximation.
-They also adopt quantization loss to optimize the selected codes towards the encoded features.
-Consequently, the VQ model can be trained in an end-to-end manner.
-Despite progress has been made, these VQ methods struggle to scale up tokenizers for better representation capacity.
 As shown in Fig.
-3, existing VQ methods simply optimize a limited number of codes towards encoded features within each backward process.
-It progressively broadens the distribution gap between the non-activated codes and the encoded features, eventually leading to codebook collapse.
-And this situation becomes more severe with the code dimension and codebook size increase.
-Rather than directly applying the straight-through estimator [1] on the selected codes, we employ this parameterization approach on the categorical distribution between visual features and all codebook embeddings to enable gradients backward to all codes.
-In such a way, the distribution between the whole codebook and the encoded features remains consistent throughout the training process.
-Consequently, IBQ achieves an extremely large codebook with high code dimension and utilization.
-3.3 Autoregressive Transformer
+3, existing VQ methods (e.g., VQ-VAE [42] and VQGAN [11]) update only a few selected codes within each backward process, progressively widening the gap between non-activated codes and encoded features, which leads to codebook collapse.
+This issue worsens as code dimension and codebook size increase.
+Instead of applying the straight-through estimator [4] on the selected codes, we employ it on the categorical distribution between visual features and all codebook embeddings to enable gradients backward to all codes.
+This promotes distribution consistency between the codebook and encoded features throughout training, allowing IBQ to scale up to extremely large codebook size with high code dimension and utilization.
+3.3 Vanilla Autoregressive Transformer
 After tokenization, the visual feature is quantized into discrete representations which are subsequently flattened in a raster-scan manner for visual generation.
-Given the discrete token index sequence $\mathcal{X}=\{x_{i}\}_{i=1}^{T}$, where $T=h^{\prime}\times w^{\prime}$, we leverage an autoregressive transformer to model the sequence dependency via next-token prediction.
+Given the discrete token index sequence $\mathcal{X}=\{x_{i}\}_{i=1}^{T}$, where $T=h^{\prime}\times w^{\prime}$, we employ an autoregressive transformer to model the sequence dependency through next-token prediction.
 Specifically, the optimization process is to maximize the log-likelihood:
@@
 Note that, since our focus is on the visual tokenizer, we adopt the vanilla architecture of autoregressive transformers akin to Llama [40] with AdaLN [28] for visual generation.
+More details can be deferred to the supplementary.
 […]
@@
 […]
-MaskBit [30] 2D 16 $\times$ 16 16 256 $\times$ 256 16,384 0 1.61 -- --
 Open-MAGVIT2 [25] 2D 16 $\times$ 16 16 256 $\times$ 256 16,384 0 1.58 0.2261 100%
@@
 Titok-S [49] 1D 128 -- $256\times 256$ 4,096 16 1.71 -- --
-Table 1: Reconstruction performance of different tokenizers on $256\times 256$ ImageNet 50k validation set. reproduced VQGAN.
+Table 1: Reconstruction performance of different tokenizers on $\boldsymbol{256\times 256}$ ImageNet 50k validation set. reproduced VQGAN.
 […]
@@
 […]
+1,024 2.24 0.2580 99%
+8,192 1.87 0.2437 98%
+16,384 1.37 0.2235 96%
+262,122 1.00 0.2030 84%
+Table 3: Impact of codebook size.
+[…]
 32 2.04 0.2408 92%
@@
 […]
-1,024 2.24 0.2580 99%
-8,192 1.87 0.2437 98%
-16,384 1.37 0.2235 96%
-262,122 1.00 0.2030 84%
-Table 4: Impact of codebook size.
+1 1.80 0.2377 99%
+2 1.55 0.2311 97%
+4 1.37 0.2235 96%
+Table 5: Benefit of larger model size.
 […]
@@
 Table 6: Performance comparison with LFQ.
-[…]
-1 1.80 0.2377 99%
-2 1.55 0.2311 97%
-4 1.37 0.2235 96%
-Table 6: Benefit of larger model size.
 Figure 4: Qualitative Reconstruction Comparison. We compare IBQ with LlamaGen and Open-MAGVIT2 tokenizer.
 4 Experiment
-Figure 5: Scaling up visual tokenizers (e.g., codebook size, code dimension and model size) improves visual soundness of reconstruction. Zoom in for a better view.
+Figure 5: Scaling up visual tokenizers (e.g., codebook size, code dimension and model size) improves visual soundness of reconstruction.
 […]
@@
 MAR-H [23] 943M 1.55 303.7 0.81 0.62
-AR VQGAN [6] 227M 18.65 80.4 0.78 0.26
+Vanilla AR VQGAN [11] 227M 18.65 80.4 0.78 0.26
 VQGAN [11] 1.4B 15.78 74.3 -- --
@@
 Open-MAGVIT2-XL [25] 1.5B 2.33 271.77 0.84 0.54
-AR IBQ-B (Ours) 342M 2.88 254.73 0.84 0.51
+Vanilla AR IBQ-B (Ours) 342M 2.88 254.73 0.84 0.51
 IBQ-L (Ours) 649M 2.45 267.48 0.83 0.52
@@
 IBQ-XXL (Ours) 2.1B 2.05 286.73 0.83 0.57
-Table 7: Class-conditional generation on $256\times 256$ ImageNet.
+Table 7: Class-conditional generation on $\boldsymbol{256\times 256}$ ImageNet.
 * specifies the generated images are $384\times 384$ and are resized to 256×256 for evaluation.
-The evaluation protocol and implementation are the same with ADM.
+The evaluation protocol and implementation are the same as ADM [10].
 4.1 Datasets and Metrics
-The training of the visual tokenizer and autoregressive transformer are both on $256\times 256$ ImageNet [4].
-For visual reconstruction, the reconstruction-FID, denoted as rFID [7], codebook utilization, and LPIPS [34] on ImageNet 50k validation set are adopted to measure the quality of reconstructed images.
-For visual generation, we measure the quality of image generation by the prevalent metrics FID, IS [21] and Precision/Recall [12].
+Both the visual tokenizers and autoregressive transformers are trained on $256\times 256$ ImageNet [9].
+For reconstruction, we measure reconstruction-FID (rFID [13]), codebook utilization, and LPIPS [50] on the ImageNet 50k validation set.
+For generation, we evaluate image quality using generation FID (gFID), Inception Score [32], and Precision/Recall [19], following ADM evaluator [10].
 4.2 Implementations Details
@@
 The visual tokenizer is trained with the following settings: an initial $41e-$ 4 learning rate with $0.01$ multi-step decay mechanism, an Adam Optimizer [18] with $\beta_{1}=0.5$, $\beta_{2}=0.9$, a total $256$ batch size with $330$ epochs, a combination of reconstruction, GAN [16], perceptual [50], commitment [11], entropy [47], double quantization losses, and LeCAM regularization [41] for training stability.
-When not specified, we use 16,384 codebook size, 256 code dimension, and 4 ResBlocks as our default tokenizer setting.
+Unless otherwise specified, we use a codebook size of 16,384, a code dimension of 256, and 4 ResBlocks as our default tokenizer setting.
 Visual Generation Setup.
-It is anticipated that the plain autoregressive visual generation method can well exhibit the effectiveness of our visual tokenizer design.
-Specifically, we adopt Llama-based architecture where the RoPE [23], SwiGLU [22], and RMSNorm [35] techniques are incorporated.
-We further introduce AdaLN [18] for better visual synthesis quality.
-The class embedding serves as the start token and the conditional of AdaLN simultaneously.
-IBQ with width $w$, depth $d$ and head $h$ follows the scaling rules proposed in [24, 25], which is formulated as:
-$w=64d,\quad h=d.$ (12)
+We use vanilla Autoregressive models ranging from 300M to 2.1B to validate the effectiveness of IBQ tokenizers in visual generation, adopting a Llama-based architecture with RoPE [36], SwiGLU [35], and RMSNorm [51].
+AdaLN [28] is also incorporated for improved visual synthesis quality.
+The class embedding serves as both the start token and AdaLN condition.
+IBQ with width $w$, depth $d$ and head $h$ follows the scaling rules proposed in [37, 39], where $w=64d,h=d$.
 All models are trained with similar settings: a base learning rate of $41e-$ 4 per $256$ batch size, an AdamW optimizer [24] with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay = $25e-$ 2, a total $768$ batch size and 300 $\sim$ 450 training epochs corresponding the model size, gradient clipping of $1.0$, and $0.1$ dropout rate for input embedding, FFN module and conditional embedding.
@@
 1 shows the quantitative reconstruction comparison between IBQ and prevalent visual tokenizers.
-It can be seen that a significant drop in codebook usage on existing VQ methods when scaling codebook size (e.g., VQGAN [6] has a 44% usage for 1024 codebook size, while a 5.9% usage for 16,384 codebook size.), and code dimension (e.g., LlamaGen [24] has a 97% usage for 8-dimension codes, while a 0.29% usage for 256-dimension codes.) Therefore, the actual representational capacity is limited by the codebook collapse.
-In contrast, our joint optimization of all codebook embedding and the visual encoder ensures the consistent distribution between them, facilitating the stable training of an large codebook size and embedding visual tokenizer with high utilization.
-Specifically, IBQ with 16,384 codebook size and 256 code dimension achieves 1.37 rFID, exceeding other VQ methods under the same downsampling rate and codebook size.
-By increasing the codebook size to 262,144, IBQ outperforms Open-MAGVIT2 [16] and achieves state-of-the-art reconstruction performance (1.00 rFID).
-We also perform a qualitative comparison with several representative VQ methods in Fig.
-4.
-IBQ exhibits better visual soundness on complex scenarios such as faces and characters.
+Existing VQ methods show a significant drop in codebook usage when scaling codebook size (e.g., VQGAN [11] has a 44% usage for 1024 codebook size, while a 5.9% usage for 16,384 codebook size.), and code dimension (e.g., LlamaGen [37] has a 97% usage for 8-dimension codes, while a 0.29% usage for 256-dimension codes.) Therefore, the actual representational capacity is limited by the codebook collapse.
+In contrast, IBQ’s joint optimization of codebook embeddings and the visual encoder maintains distribution consistency, enabling stable training of large-scale codebook with high utilization.
+Specifically, IBQ with 16,384 codebook size and 256 code dimension achieves 1.37 rFID, outperforming other VQ methods at the same downsampling rate and codebook size.
+Increasing the codebook size to 262,144, IBQ achieves state-of-the-art reconstruction with 1.00 rFID, surpassing Open-MAGVIT2 [25].
+A qualitative comparison in Fig.
+4 shows IBQ’s superior visual quality in complex scenarios such as faces and characters.
+Note that, we observe that incorporating additional facial data yields consistent improvements (see supplementary materials).
 Visual Generation.
 In Tab.
-7, we compare IBQ with other generative models, including Diffusion models, AR models, and variants of AR models (VAR [25] and MAR [14]) on class-conditional image generation task.
-With the powerful visual tokenizers of IBQ, our models show consistent improvements when scaling up the model size (from 300M to 2.1B), and outperform all previous vanilla autoregressive models at different scales of model size.
+7, we compare IBQ with other generative models, including Diffusion models, AR models, and variants of AR models (VAR [39] and MAR [23]) on class-conditional image generation.
+With powerful IBQ tokenizers, our models show consistent improvements when scaling up the model size (from 300M to 2.1B), and outperform all previous vanilla autoregressive models at different scales of model size.
 Moreover, IBQ outperforms the diffusion-based model DiT [28], and achieves comparable results with the variants of AR models.
@@
 Therefore, we believe that with our stronger tokenizers, the AR models and their variants can be boosted further.
-4.4 Scaling Up IBQ
-[…]
-IBQ shows promising scaling capacity in three aspects: 1) Codebook Size: As shown in Tab.
-6, there is a significant improvement in reconstruction quality as the codebook size enlarges from 1024 to 16,384.
-Furthermore, IBQ enables high codebook utilization and consistent gain in visual soundness even training with 262,144 codes.
-2) Model Size: Tab.
-6 reveals that by extending the number of ResBlock both in the encoder and the decoder, the boost in reconstruction performance can be guaranteed.
-3) Code Dimension: interestingly, we observe significant codebook usage increasements when scaling up code dimension.
-We assume that low-dimensional codes are less discriminative where similar ones tend to be clustered.
-It indicates that the representative codes are more likely to be selected under our global updating strategy.
-In contrast, codes with high-dimensional embedding are highly informative since they are mutually sparse in the representation space.
-Therefore, these codes can be evenly selected during training, which ensures high utilization with better performance.
-With the above factors, we realize a super large codebook of 262,144 codebook size and 256 dimensions with high codebook usage (84%), achieving the state-of-the-art reconstruction performance (1.00 rFID).
+4.4 Scaling Up IBQ Tokenizers
+Existing VQ methods struggle to scale up due to the codebook collapse.
+For example, LlamaGen [37] sees a significant drop in usage and rFID when increasing the code dimension from 8 to 256 (97% $\rightarrow$ 0.29%, 2.19 rFID $\rightarrow$ 9.21 rFID), as shown in Tab.
+1.
+This is due to their partial updates during training, which progressively widens the distribution gap between non-activated codes and encoded features.
+Scaling Up Tokenizers Improves Reconstruction.
+IBQ tokenizers show promising scaling capacity for reconstruction in three aspects: 1) Codebook Size: As shown in Tab.
+6, reconstruction quality improves significantly as the codebook size increases from 1,024 to 16,384, with high utilization and consistent visual soundness even at 262,144 codes.
+2) Code Dimension: interestingly, we observe a notable increase in codebook usage when scaling code dimension in Tab.
+6.
+We assume that low-dimensional codes are less discriminative and tend to be clustered, indicating that representative codes are more likely to be selected under our global updating strategy.
+In contrast, high-dimensional codes are highly informative due to their sparsity in the representation space, allowing for more even selection during training, which ensures high utilization with better performance.
+3) Model Size: Tab.
+6 reveals that by increasing the number of ResBlock both in both the encoder and decoder, a higher reconstruction performance can be guaranteed.
+With these factors, we realize a super large codebook of 262,144 codebook size and 256 dimensions with high codebook usage (84%), achieving the state-of-the-art reconstruction performance (1.00 rFID).
 To better illustrate the scaling properties, we also provide visualizations in Fig.
 5.
+Figure 6: Scaling up IBQ tokenizers enables better generation, especially with larger autoregressive models (e.g., 1B param.).
+Scaling Up Tokenizers Improves Generation.
+Scaling up IBQ tokenizers also enhances generation quality.
+As shown in Fig.
+6, increasing the codebook size significantly improves reconstruction and generation FID, with a similar trend observed when scaling code dimensions.
+Moreover, with larger autoregressive models (e.g., 1B parameters), the improvement in generation quality becomes more remarkable, suggesting that scaling up generative models can further unlock the potential of IBQ tokenizers.
 4.5 Ablation Studies
@@
 For fair comparisons, we adopt LFQ [25] with 16,384 codes and replace its asymmetric token factorization with our vanilla transformer architecture.
-We compare with LFQ on both reconstruction and generation in Tab.
-6, and our proposed IBQ performs better, which demonstrates increasing code dimension can improve the reconstruction ability of the visual tokenizer and further boost the visual generation.
+Tab.
+6 shows that IBQ outperforms LFQ in both reconstruction and generation, which demonstrates increasing code dimension can improve the reconstruction ability of the visual tokenizer and further boost the visual generation.
 5 Conclusion
-In this paper, we identify the bottleneck in scaling tokenizers arising from the partial-update strategy in current VQ methods, which progressively broadens the distribution gap between encoded features and non-activated codes, ultimately leading to codebook collapse.
-To tackle this challenge, we propose a simple yet effective vector quantization method, coined as Index Backpropagation Quantization (IBQ), for scalable tokenizer training, which updates all codes by applying the straight-through estimator on the categorical distribution between visual features and all codebook embeddings, thereby maintaining consistent distribution between the entire codebook and encoded features.
-Experiments on ImageNet demonstrate that IBQ enables a high-utilization, large-scale visual tokenizer with improved performance in both reconstruction (1.00 rFID) and generation (2.05 gFID), confirming the scalability and effectiveness of our method.
+In this paper, we identify the bottleneck in scaling tokenizers (e.g., codebook size), stemming from the partial-update strategy in current VQ methods, which progressively enlarge the distribution gap between encoded features and non-activated codes, eventually leading to codebook collapse.
+To address this challenge, we propose a simple yet effective vector quantization method, termed as Index Backpropagation Quantization (IBQ), for scalable tokenizer training, which updates all codes by applying the straight-through estimator on the categorical distribution over visual features and all codebook embeddings, thereby maintaining consistent distribution between the entire codebook and encoded features.
+Experiments on ImageNet demonstrate that IBQ enables a high-utilization, large-scale visual tokenizer with improved performance in both reconstruction (1.00 rFID) and generation (2.05 gFID).
 Appendix
-Appendix A Autoregressive Model Configurations
-We show the detailed autoregressive model configurations in Tab. 8. We scale up the AR models from 342M to 2.1B following the scaling rules proposed in VAR [25].
-Appendix B Comparison with Soft Vector Quantization
-To comprehensively illustrate the rationality of our IBQ, we compare it with another global update method, Soft Vector Quantization dubbed as Soft VQ.
-During training, it adopts the weighted average of all code embeddings as the quantized feature $v_{q}$ and incorporates a cosine decay schedule of the temperature ranging from $0.9$ to $1e$-$6$ for one-hot vector approximation.
-As for inference, it switches back to the original VQGAN way, which selects the code with the highest probability for hard quantization.
-As shown in Tab.
-9, Soft VQ is far behind IBQ in both reconstruction quality and codebook usage.
-In the experiments, we observe that the training process of Soft VQ corrupts within a few epochs ($<10$).
-This may stem from the unstable adversarial training where the adaptive weight of the GAN loss appears enormous and ends up with NAN.
-In addition, the soft-to-hard manner for one-hot vector approximation brings more difficulty in optimization and incurs inconsistency of quantization between training and inference, as demonstrated by a significant reconstruction quality drop ($16.17\text{rFID}\rightarrow 233.17\text{rFID}$).
-Moreover, we provide an in-depth investigation by visualizing the distribution between the codebook and encoded features of Soft VQ.
-As shown in Fig.
-6, although all-code updating strategy is enabled, the inappropriate quantization process tends to cluster codes mistakenly which subsequently suffers from low codebook usage (2.5%).
-We speculate that the force of the weighted average of code embeddings toward the encoded feature will smooth the codebook representation and result in similar and less informative code embeddings.
-In contrast, IBQ adopts hard quantization with index backpropagation.
-The hard quantization only involves the selected codes toward the encoded features for discriminative representation, thus ensuring precise quantization, while index backpropagation performs joint optimization of all codebook and the visual encoder to achieve consistent distribution.
-Considering the factors above, our proposed IBQ shows dominance in both reconstruction quality and codebook utilization.
-Appendix C Additional Visualizations
-We provide more qualitative reconstruction and generation samples in Fig. 7 and Fig. 8, respectively.
-Appendix D Limitation and Future Work
-In this work, we mainly focus on visual tokenizers through scalable training, and adopt vanilla autoregressive transformers for visual generation.
-We believe that combining our powerful visual tokenizers with advanced AR models is a promising way to improve autoregressive visual generation.
-In addition, our models are trained on ImageNet [4], limiting the generalization ability.
-Pretraining tokenizers on more data and training AR models on text-image pairs would be helpful.
-We leave these for future work.
-Figure 6: Distribution Gap. The T-SNE results of the codebook (16,384 codebook size and 256 dimension) and sampled encoded features.
-Model Parameters Width $w$ Head $h$ Depth $d$
-IBQ-B 342M 16 16 1024
-IBQ-L 649M 20 20 1280
-IBQ-XL 1.1B 24 24 1536
-IBQ-XXL 2.1B 30 30 1920
+Model Parameters Width $w$ Head $h$ Depth $d$ Lr Batch Size Epoch
+IBQ-B 342M 16 16 1024 3e-4 768 300
+IBQ-L 649M 20 20 1280 3e-4 768 350
+IBQ-XL 1.1B 24 24 1536 3e-4 768 400
+IBQ-XXL 2.1B 30 30 1920 3e-4 768 450
 Table 8: Model sizes and architecture configurations of IBQ.
@@
 denotes IBQ with the same training epochs as Soft VQ.
-Figure 7: The upper part illustrates the IBQ tokenizer tested at 1024 $\times$ 1024 Unsplash.
+Model Codebook Size Parameters Memory Time/epoch Usage
+VQGAN 1,024 89.6M 19.5G 3h15min 44%
+8,192 91.5M 19.7G 3h18min -
+16,384 93.6M 19.8G 3h21min 5.3%
+262,144 156M 21.2G 4h $\sim$0%
+IBQ 1,024 89.6M 19.5G 3h20min 99%
+8,192 91.5M 19.7G 3h30min 98%
+16,384 93.6M 20G 3h40min 96%
+262,144 156M 30.5G 9h 84%
+Table 10: Training computational costs comparison between VQGAN and IBQ. (Tested on 8 A6000 gpus)
+Appendix A Autoregressive Model Configurations
+We show the detailed autoregressive model configurations and training settings in Tab.
+8.
+We scale up the autoregressive models from 300M to 2.1B parameters, following the scaling rules proposed in VAR [39].
+Appendix B Comparison with Soft Vector Quantization
+To comprehensively illustrate the rationality of our IBQ, we compare it with another global update method, Soft Vector Quantization (Soft VQ).
+During training, it adopts the weighted average of all code embeddings as the quantized feature $v_{q}$ and incorporates a cosine decay schedule of the temperature ranging from $0.9$ to $1e$-$6$ for one-hot vector approximation.
+As for inference, it switches back to the original VQGAN way, which selects the code with the highest probability for hard quantization.
+As shown in Tab.
+9, Soft VQ is far behind IBQ in both reconstruction quality and codebook usage.
+In the experiments, we observe that the training process of Soft VQ corrupts within a few epochs ($<10$).
+This may stem from the unstable adversarial training where the adaptive weight of the GAN loss appears enormous and ends up with NAN.
+In addition, the soft-to-hard manner for one-hot vector approximation brings more difficulty in optimization and incurs inconsistency of quantization between training and inference, as demonstrated by a significant reconstruction quality drop ($16.17\text{rFID}\rightarrow 233.17\text{rFID}$).
+Moreover, we provide an in-depth investigation by visualizing the distribution between the codebook and encoded features of Soft VQ.
+As shown in Fig.
+7, although all-code updating strategy is enabled, the inappropriate quantization process tends to cluster codes mistakenly, resulting in low codebook usage (2.5%).
+We speculate that the force of the weighted average of code embeddings toward the encoded feature will smooth the codebook representation and result in similar and less informative code embeddings.
+In contrast, IBQ adopts hard quantization with index backpropagation.
+The hard quantization only involves the selected codes toward the encoded features for discriminative representation, thus ensuring precise quantization, while index backpropagation performs joint optimization of the entire codebook and visual encoder to achieve consistent distribution.
+Considering the factors above, our proposed IBQ shows dominance in both reconstruction quality and codebook utilization.
+Figure 7: Distribution Gap. The T-SNE results of the codebook (16,384 codebook size and 256 dimension) and sampled encoded features.
+Figure 8: Face reconstruction comparison. Scaling up tokenizers and finetuning tokenizers on face data can effectively improve facial reconstruction performance.
+Method Ratio Codebook MS-COCO 2017 Imagenet-1k
+[…]
+LlamaGen 16 16384 8.40 20.28 0.55 2.47 20.65 0.54
+Show-o 16 8192 9.26 20.90 0.59 3.50 21.34 0.59
+Cosmos 16 64000 11.97 19.22 0.48 4.57 19.93 0.49
+Open-MAGVIT2 16 16384 7.93 22.21 0.62 2.55 22.21 0.62
+Open-MAGVIT2 16 262144 6.76 22.31 0.65 1.67 22.70 0.64
+IBQ (Ours) 16 16384 7.67 21.58 0.62 2.06 22.01 0.61
+IBQ (Ours) 16 262144 6.79 22.28 0.65 1.53 22.69 0.64
+Table 11: Zero-shot reconstruction performance on ImageNet 50k validation set and MS-COCO val2017.
+The tokenizers are trained with large-scale general-domain datasets and aim to serve text-conditional image generation.
+The results are reported under the same setup for fair comparison (text in gray signifies the results directly from Cosmos report).
+$\dagger$ indicates that LlamaGen loads the model initially trained on Imagenet while the others are training from scratch, i.e., MS-COCO and Imagenet-1k are excluded from training data.
+Appendix C Training Costs
+We evaluate the training costs of VQGAN and IBQ under varying codebook sizes using 8 A6000 GPUs.
+As shown in Tab.
+10, the all-codes updating mechanism of IBQ incurs only a marginal increase in training costs compared to VQGAN when the codebook size is up to 16,384, yet it significantly improves codebook utilization.
+Specifically, IBQ introduces an additional 0.2 GB of memory usage and extends training time by 19 minutes, but increases codebook utilization from 5.3% to 96%.
+Furthermore, VQGAN fails to train with an extremely large codebook (i.e., 262,144 entries), whereas IBQ successfully achieves 84% utilization.
+Appendix D Pretraining Tokenizer
+We further unveil the representation capacity of our tokenizer by pretraining IBQ on large-scale domain datasets, i.e., 1) General: CapFusion [48], LAION-COCO [20], CC12M [8] and CC3M [34].
+2) High-quality: LAION-aesthetics-12M [^1] https://huggingface.co/datasets/dclure/laion-aesthetics-12m-umap, LAION-aesthetics [33], JourneyDB [27] and LAION-HD [^2] https://huggingface.co/datasets/yuvalkirstain/laion-hd-subset.
+We follow the same training settings stated in the manuscript while the training steps are $\sim$ 800,000.
+It can be seen in the Tab.
+11 that IBQ achieves state-of-the-art performance compared to concurrent methods such as Cosmos [1], Show-o [45].
+Although some recent efforts in residual tokenization [29, 12] can achieve better results, they are not listed here because residual techniques are orthogonal and compatible with IBQ.
+It is anticipated that our improvement on the naive quantization method better benefits the unified visual understanding and generation models compared to the residual one.
+Appendix E Improving Face Reconstrution
+Visual tokenizers trained on ImageNet may not perform as expected for face reconstruction.
+Increasing the codebook size can effectively mitigate this limitation.
+As shown in Fig.
+8, increasing the codebook size from 16,384 to 262,144 leads to improved face reconstruction quality.
+Additionally, incorporating face data into the training set or fine-tuning on face-specific datasets are effective strategies for further enhancement.
+In particular, fine-tuning IBQ on the FFHQ dataset further enhances reconstruction performance.
+Appendix F Additional Visualizations
+We provide more qualitative reconstruction and generation samples in Fig. 9 and Fig. 10, respectively.
+Figure 9: Reconstruction samples.
+The upper part illustrates the IBQ tokenizer tested at 1024 $\times$ 1024 Unsplash.
 While the second part showcases the IBQ tokenizer tested at 256 $\times$ 256 Imagenet.
 (a) indicates the original images and (b) signifies the reconstructions.
-Figure 8: We showcases the 256 $\times$ 256 class conditional generation samples on Imagenet.
+Figure 10: Generation samples. We showcase the 256 $\times$ 256 class conditional generation samples on Imagenet.